# Spatial-Temporal-Spectral Unified Modeling for Remote Sensing Dense Prediction

### v2

@@ III-D Local-Global Window Attention
Global attention mechanisms have become widely adopted in remote sensing tasks due to their strong capability to model long-range dependencies. However, their quadratic computational complexity with respect to sequence length poses a major challenge when dealing with large-scale remote sensing data. Moreover, global attention primarily captures holistic contextual features, often neglecting local spatial patterns, which are equally important for dense prediction tasks. To address this issue, this study proposes the Local-Global Window Attention module (LGWA), which improves both computational efficiency and modeling capability by jointly capturing fine-grained local features and global contextual information.
Figure 4: Architecture of the proposed LGWA-based Basic block, employing multiple local window to extract local feature and single global window to extract global feature.
LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 4. These configurations determine the receptive field and sensitivity of the attention operation within each window and directly influence computational cost. Selecting appropriate window configurations in different layers allows for an optimal trade-off between feature richness and resource efficiency. For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@
Through this modular design, LGWA effectively captures both local structural details and global semantic context. The flexible design of window configurations allows the module to balance modeling power and computational efficiency, making it well-suited for multi-scale, high-resolution remote sensing applications.
@@ IV Experiments and Results
In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets. These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatio-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets. The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal semantic segmentation, binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
@@
Dynamic Earthnet [41] LULC SCD &SS &BCD 24 24 4 6 1024×1024 3 54750
@@ IV-B Baseline
To evaluate the effectiveness of the proposed STSUN, we conducted comparative experiments with various benchmark methods on building and LULC scene datasets. Since STSUN is adaptable to multiple datasets, it was trained and validated across all datasets in various scenarios, whereas the benchmark methods were trained and validated on individual datasets.
@@
On the two LULC scene datasets, the compared CNN-based models include U-Net [44], Deepabv3+ [48], DANet [71], ResUNet-a [49], DASSN [72], HCANet [73], RAANet [74], A2-FPN [75] and CAC [76], the compared transformer-based models include LANet [77], SCAttNet [78], and TSViT [79], the CNN-transformer hybrid models include SAPNet [80], SSCBNet [81], UTAE [82], A2Net [83], SCanNet [84] and TSSCD [85].
@@ IV-C Implementation Details
3) Evaluation Metrics: The performance of the proposed models was quantitatively assessed using five principal metrics: overall accuracy (OA), precision (P), recall (R), F1-score, and intersection over union (IoU). For multi-temporal tasks and multi-category tasks, the average F1-score (AF) and mean IoU (mIoU) will be used. In addition, following the settings of the DynamicEarthnet dataset [41], we use the semantic change segmentation (SCS) metric, classagnostic binary change score (BC) and semantic segmentation score among changed pixels (SC) metrics to evaluate the model’s performance on this dataset. Precision provides a measure of the incidence of false positives, while recall reflects the occurrence of false negatives. Due to the inherently inverse relationship between these two metrics, simultaneously securing high scores for both represents a significant challenge. The F1-score, calculated as the harmonic mean of precision and recall, offers a balanced assessment. The IoU metric evaluates spatial accuracy by quantifying the overlap between the predicted and ground truth changed pixels relative to their union.
IV-D Ablation Study
To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset. The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions. As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”). Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension. The DUM is then used to map the combined channel dimension into a unified feature. In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task. Concurrently, LGWA was substituted with standard global attention. This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
TABLE II: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
Strategy Attention F1(%) IoU(%) OA(%)
Coupled Unification Global Attention 62.87 45.85 94.38
Coupled Unification LGWA 63.22 46.22 94.72
Decoupled Unification Global Attention 65.91 49.15 97.21
Decoupled Unification LGWA 66.48 49.79 98.05
Table II displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives. Furthermore, their combined use leads to even greater performance improvements.
IV-E Overall Comparison
IV-E1 Building Scenario
The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection. STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
@@ IV-E1 Building Scenario
On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed STSUN achieves state-of-the-art performance. As detailed in Table III, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%. This superior performance can be attributed to STSUN, particularly its spatial unification component, which effectively processes varying image size, and the Local-Global Window Attention mechanism. The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
TABLE III: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
@@
For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table IV. The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points. STSUN excels here due to its inherent design for unified temporal and spectral modeling. The STSUN component allows the network to effectively learn temporal evolutionary patterns and relationships between buildings and other objects, leading to more precise change maps.
TABLE IV: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
@@
In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, STSUN surpasses compared methods. Table V shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%. The strength of STSUN on this dataset lies in its robust temporal modeling capabilities, which allows for consistent feature representation across different time points. Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
TABLE V: Accuracy Comparison on the LEVIR-CD Dataset. The Best Values Are Highlighted in Bold.
@@
Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%. The TSCD dataset requires robust modeling of temporal dependencies across multiple observations. STSUN is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of buildings, which enables STSUN to accurately identify changes in complex, evolving landscapes.
TABLE VI: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
@@
Figure 5 presents inference results of STSUN from all four datasets. Visually, STSUN consistently produces accurate and complete segmentation and change maps. The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations. For instance, in building segmentation on the WHU dataset (Figure 5(a)), STSUN generates sharp edges and complete building shapes. Similarly, for change detection tasks on WHU-CD (Figure 5(b)), LEVIR-CD (Figure 5(c)), and TSCD (Figure 5(d)), STSUN demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas. This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of STSUN and LGWA, to learn highly discriminative features and effectively model contextual relationships across the spatial, temporal and spectral dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
IV-E2 LULC Scenario
@@
On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII. This leading performance can be attributed to STSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information. This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery.
TABLE VII: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
@@
For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VIII. The success of STSUN on this dataset underscores the effectiveness of the STSUN in explicitly modeling the temporal dimension. By leveraging temporal metadata to ensure consistency across sequences of varying lengths, STSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
TABLE VIII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.

### v3

@@ III-D Local-Global Window Attention
After unifying the diverse spatial-temporal-spectral input and output data, STSUN requires a powerful feature extraction capability to be effectively applied to various remote sensing scenarios and tasks. Therefore, this study proposes the Local-Global Window Attention module. This module performs self-attention operations within multiple local windows of varying shapes and a single global window, As illustrated in Figure 4. This design enables the joint capture of fine-grained local details and coarse-grained global features, thereby effectively extracting the local characteristics of various ground objects and modeling the global contextual relationships among them, which enhances the model’s robustness across diverse remote sensing applications.
The LGWA module incorporates three local window attention mechanisms with varying sizes and configurations, alongside a single global attention mechanism. The distinct shapes of the local windows are specifically designed to extract different levels of local features. The horizontal window concentrates on capturing horizontally-oriented local details, the vertical window focuses on vertically-aligned local features, and the rectangular window is dedicated to extracting omnidirectional local information. This multi-faceted approach ensures the comprehensive extraction of multi-level features from ground objects. Subsequently, the global attention mechanism models the global contextual relationships among these objects based on their extracted local features, which enhances the overall performance of the model.
Figure 4: Architecture of the LGWA-based Local Global Block, employing multiple local windows to extract local feature and single global window to extract global feature.
For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@ IV Experiments and Results
In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets. These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatial-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets. The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
@@
Dynamic Earthnet [41] LULC BCD &SCD 24 24 4 6 1024×1024 3 54750
@@ IV-B Baseline
To evaluate the effectiveness of the proposed STSUN, we conducted comparative experiments with various benchmark methods on building and LULC scene datasets. Since STSUN is adaptable to multiple datasets, it was trained across all datasets in each scenario, whereas the benchmark methods were trained on individual datasets.
@@
On the two LULC scene datasets, the compared CNN-based models include U-Net [44], Deepabv3+ [48], DANet [71], ResUNet-a [49], DASSN [72], HCANet [73], RAANet [74], MSAFNet [75], A2-FPN [76] and CAC [77], the compared transformer-based models include LANet [78], SCAttNet [79], CLCFormer [80] and TSViT [81], the CNN-transformer hybrid models include SAPNet [82], SSCBNet [83], UTAE [84], A2Net [85], SCanNet [86] and TSSCD [87].
@@ IV-C Implementation Details
3) Evaluation Metrics: The performance of the proposed models was quantitatively assessed using five principal metrics: overall accuracy (OA), precision (P), recall (R), F1-score, and intersection over union (IoU). For multi-temporal tasks and multi-category tasks, the average F1-score (AF) and mean IoU (mIoU) will be used. In addition, following the settings of the DynamicEarthnet dataset [41], we use the semantic change segmentation (SCS) metric, classagnostic binary change score (BC) and semantic segmentation score among changed pixels (SC) metrics to evaluate the model’s performance on this dataset.
IV-D Overall Comparison on the Building Scenario
The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection. STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table II, III, IV and V.
@@ IV-D Overall Comparison on the Building Scenario
On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed STSUN achieves state-of-the-art performance. As detailed in Table II, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%. This superior performance can be attributed to STSUN, particularly its spatial unification component, which effectively processes varying image size, and the Local-Global Window Attention mechanism. The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
TABLE II: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
@@
For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table III. The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points. STSUN excels here due to its inherent design for unified temporal and spectral modeling. The STSUN component allows the network to effectively learn temporal evolutionary patterns and relationships between buildings and other objects, leading to more precise change maps.
TABLE III: Accuracy Comparison on the WHU-CD Dataset. The Best Values Are Highlighted in Bold.
@@
In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, STSUN surpasses compared methods. Table IV shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%. The strength of STSUN on this dataset lies in its robust temporal modeling capabilities, which allows for consistent feature representation across different time points. Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
TABLE IV: Accuracy Comparison on the LEVIR-CD Dataset. The Best Values Are Highlighted in Bold.
@@
Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table V, with an F1-score of 66.48% and an IoU of 49.79%. The TSCD dataset requires robust modeling of temporal dependencies across multiple observations. STSUN is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of buildings, which enables STSUN to accurately identify changes in complex, evolving landscapes.
TABLE V: Accuracy Comparison on the TSCD Dataset. The Best Values Are Highlighted in Bold.
@@
Figure 5 presents inference results of STSUN from all four datasets. Visually, STSUN consistently produces accurate and complete segmentation and change maps. The results exhibit few false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations. For instance, in building segmentation on the WHU dataset (Figure 5(a)), STSUN generates sharp edges and complete building shapes. Similarly, for change detection tasks on WHU-CD (Figure 5(b)), LEVIR-CD (Figure 5(c)), and TSCD (Figure 5(d)), STSUN demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas. This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of STSUN and LGWA, to learn highly discriminative features and effectively model contextual relationships across the spatial, temporal and spectral dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
IV-E Overall Comparison on the LULC Scenario
@@
On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VI. This leading performance can be attributed to STSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information. This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery.
TABLE VI: Accuracy Comparison on the LoveDA Dataset. The Best Values Are Highlighted in Bold.
@@
For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VII. The success of STSUN on this dataset underscores the effectiveness of the STSUN in explicitly modeling the temporal dimension. By leveraging temporal metadata to ensure consistency across sequences of varying lengths, STSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
TABLE VII: Accuracy Comparison on the DynamicEarthnet Dataset. The Best Values Are Highlighted in Bold.
@@
IV-F Ablation Study
To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset. The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions. As the temporal dimension is strongly correlated with dense prediction task types, this strategy enables joint modeling of multiple tasks on independent dimensions, learning the complementarity between each task. As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”). Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension. The DUM is then used to map the combined channel dimension into a unified feature. In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task. Concurrently, LGWA was substituted with standard global attention. This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
TABLE VIII: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
Strategy Attention F1(%) IoU(%) OA(%)
Coupled Unification Global Attention 62.87 45.85 94.38
Coupled Unification LGWA 63.22 46.22 94.72
Decoupled Unification Global Attention 65.91 49.15 97.21
Decoupled Unification LGWA 66.48 49.79 98.05
Table VIII displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives. Furthermore, their combined use leads to even greater performance improvements.
IV-G Unification of Dense Prediction Tasks
Semantic segmentation, binary change detection, and semantic change detection are prevalent dense prediction tasks in remote sensing, exhibiting significant similarities and complementarities. To validate the capability of STSUN to unify these diverse dense prediction tasks, we conducted a comparison experiment on four datasets in the building scenario. Specifically, we compare the performance of STSUN models trained individually on each of the four datasets against a single STSUN model trained jointly on all datasets. For a given dataset, the former is denoted as $STSUN_{\{dataset\}\_single}$ and the latter as $STSUN_{\{dataset\}\_unified}$.
The results, presented in Table IX, demonstrate that $STSUN_{\{dataset\}\_unified}$ consistently outperforms its single-task counterparts across all building datasets. This performance gain is attributed to two factors. First, joint training on data from multiple dense prediction tasks exposes the model to a richer and more diverse set of samples, enhancing its ability to learn comprehensive spatial-temporal-spectral features from remote sensing imagery. Second, the model capitalizes on the complementary nature of these tasks. This synergy, analogous to a multi-task learning paradigm, enables the learning of more powerful and robust feature representations, thereby elevating the model’s performance across all individual tasks.
TABLE IX: Comparison over Task Unification on Four Building Scenario Datasets. The Best Values Are Highlighted in Bold.
Methods P (%) R (%) F1 (%) IoU (%)
$STSUN_{whu\_single}$ 95.24 94.30 94.77 90.06
$STSUN_{whu\_unified}$ 95.71 94.87 95.29 91.00
$STSUN_{whucd\_single}$ 92.22 90.89 91.55 84.42
$STSUN_{whucd\_unified}$ 93.07 91.20 92.13 85.40
$STSUN_{levircd\_single}$ 92.94 89.63 91.25 83.92
$STSUN_{levircd\_unified}$ 93.17 90.07 91.59 84.49
$STSUN_{tscd\_single}$ 66.91 65.08 65.98 49.23
$STSUN_{tscd\_unified}$ 67.32 65.67 66.48 49.79
IV-H Dense Prediction with Flexible Semantic Category Set
Remote sensing scenarios often encompass distinct sets of ground objects, necessitating different sets of semantic categories. For instance, the two LULC datasets employed in this study exhibit this variance: the LoveDA dataset uses the semantic set {building, road, water, barren, forest, agriculture, background}, whereas the DynamicEarthNet dataset uses {impervious surface, agriculture, forest, wetlands, soil, water}. To demonstrate STSUN’s ability to handle flexible semantic category sets for dense prediction, we designed an experiment using these two datasets. For this experiment, the DynamicEarthNet dataset was used exclusively for semantic segmentation, utilizing only the image pairs with semantic annotations. We compare four model configurations:
  1. 1.
    $STSUN_{loveda}$, trained exclusively on the LoveDA dataset to predict its 7-category set.
  2. 2.
    $STSUN_{dynamic.}$, trained exclusively on the DynamicEarthNet dataset to predict its 6-category set.
  3. 3.
    $STSUN_{fixed}$, trained jointly on both datasets to predict a fixed, 10-category set corresponding to the union of their individual category sets.
  4. 4.
    $STSUN_{flexible}$, trained jointly on both datasets but dynamically predicting from the appropriate category subset for each respective dataset.
As shown in Table X, $STSUN_{flexible}$ outperforms $STSUN_{fixed}$ on both datasets, while achieving performance comparable to the specialist models, $STSUN_{loveda}$ and $STSUN_{dynamic.}$. The degraded performance of $STSUN_{fixed}$ is expected. This model is constrained to predict over the union of all categories, even for an image from a scene that does not contain certain categories. Forcing a single, fixed-size output space across scenes with disparate semantic sets introduces ambiguity and negatively impacts performance on each respective scene. In contrast, $STSUN_{flexible}$ dynamically adapts its predictive output to the relevant subset of semantic categories for each scene, which explains why its performance is on par with the individually trained models. Crucially, however, $STSUN_{flexible}$ offers superior model efficiency, as a single, unified model can be deployed across scenes with different semantic category sets without requiring any additional training or fine-tuning.
TABLE X: Comparison over Category Unification on Two LULC Scenario Datasets. The Best Values Are Highlighted in Bold.
$Model_{loveda}$ AF(%) OA(%) mIoU(%)
$STSUN_{loveda}$ 74.66 71.70 65.59
$STSUN_{fixed}$ 74.03 70.49 64.67
$STSUN_{flexible}$ 74.81 71.82 65.73
$Model_{dynamic.}$ SCS BC SC mIoU(%)
$STSUN_{dynamic.}$ 30 38.4 21.6 55.3
$STSUN_{fixed}$ 29.4 38.7 20.1 54.2
$STSUN_{flexible}$ 30.3 38.9 21.8 55.7